# Evaluating Test-Time Scaling LLMs for Legal Reasoning: OpenAI o1, DeepSeek-R1, and Beyond

### v1

@@ Abstract
Recently, Test-Time Scaling Large Language Models (LLMs), such as DeepSeek-R1 and OpenAI o1, have demonstrated exceptional capabilities across various domains and tasks, particularly in reasoning. While these models have shown impressive performance on general language tasks, their effectiveness in specialized fields like legal remains unclear. To address this, we present a preliminary evaluation of LLMs in various legal scenarios, covering both Chinese and English legal tasks. Our analysis includes 9 LLMs and 17 legal tasks, with a focus on newly published and more complex challenges such as multi-defendant legal judgments and legal argument reasoning. Our findings indicate that, despite DeepSeek-R1 and OpenAI o1 being among the most powerful models, their legal reasoning capabilities are still lacking. Specifically, these models score below 80% on seven Chinese legal reasoning tasks and below 80% on two English legal reasoning tasks. This suggests that, even among the most advanced reasoning models, legal reasoning abilities remain underdeveloped.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have markedly reduced the gap toward Artificial General Intelligence (AGI)  Meta (2024); Team (2024); Openai (2024a, b); team (2025); Anthropic (2025). Exemplary models such as OpenAI GPT-4o  Openai (2024a), Qwen2.5  Team (2024), and Meta Llama-3.1  Meta (2024) highlight this progress. Among them, OpenAI has introduced a series of models, referred as o1  Openai (2024b), which improve the reasoning performance across diverse domains including mathematics, coding, and scientific inquiry by increasing the test-time computation through extending the length of the chain-of-thought. However, o1 are proprietary models, and their technical details and weights have not been disclosed. Following OpenAI o1, several studies have been explored to match their comprehensive reasoning capabilities, such as the Process Reward Model (PRM)  Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2024), Reinforcement Learning(RL) techniques  Kumar et al. (2024), and search algorithms like Monte Carlo Tree Search and Beam Search  Feng et al. (2024); Trinh et al. (2024). However, the performance is still lacking.
Figure 1: Overall results of the LLMs in Chinese and English legal tasks. We show performance on all legal datasets, excluding Sara_N for comparative analysis. DeepSeek-R1 and o1-preview demonstrate a clear performance advantage over the General and other TTS LLMs.
More recently, DeepSeek has open-sourced its test-time compute LLM, DeepSeek-R1 DeepSeek-AI et al. (2025). DeepSeek-AI et al. (2025) found that large-scale reinforcement learning (RL) with simple rule-based reward can already incentive the reasoning ability of LLMs, e.g., the occurrence of an "aha moment", where the model is referred to as Deekseek-R1-Zero. However, only training with RL leads to poor readability and language mixing issues. To address these challenges, a four-stage pipeline is proposed, comprising: (i) a cold-start stage, (ii) reasoning-oriented reinforcement learning, (iii) rejection sampling and supervised fine-tuning, and (iv) reinforcement learning across all scenarios. Upon completing these stages, the resulting model, DeepSeek-R1, effectively mitigates the limitations of pure reinforcement learning while demonstrating strong reasoning capabilities and competitive performance across a broad spectrum of general-purpose tasks. Notably, DeepSeek-R1 outperforms OpenAI o1-1217 on several reasoning benchmarks, attaining an accuracy of 79.8% and 97.3% on AIME 2024 and MATH-500, respectively. Concurrently, other test-time scaling LLMs have emerged, including Qwen Team’s QWQ-32B-Preview  Qwen Team (2024) and Zhipu AI’s GLM-zero-preview  [^1] https://bigmodel.cn/dev/api/normal-model/glm-zero-preview.
Despite significant progress in enhancing LLMs’ reasoning capabilities, questions remain about their performance in legal reasoning tasks, which require not only relevant legal knowledge but also the ability to apply that knowledge through reasoning. Do models that perform well in general reasoning exhibit comparable proficiency in the legal domain? If not, what are the underlying reasons for this discrepancy? Prior research has explored GPT-4’s performance across various legal tasks, including comparisons with earlier GPT models on legal text annotation  Savelka and Ashley (2023), generating factual explanations of legislative terms  Savelka et al. (2023), and conducting thematic analysis in empirical legal studies  Drápal et al. (2023). However, none of these studies offer a comprehensive and in-depth analysis of LLMs’ legal reasoning capabilities.
To address this gap, we conduct a comprehensive evaluation of 17 legal reasoning tasks (7 in English and 10 in Chinese) from prior research, testing a range of representative LLMs. Our assessment includes both test-time scaling LLMs and general-purpose LLMs, aiming to systematically analyze their performance in legal reasoning tasks. Key findings from our evaluation include:
@@
    In the Chinese legal reasoning tasks, the DeepSeek-R1 model demonstrates the highest performance among the evaluated models, with OpenAI o1-preview following closely behind. However, all evaluated LLMs encounter significant challenges when dealing with tasks that demand complex reasoning capabilities, such as controversial focus extraction, judicial examination, and multi-defendant legal judgment prediction.
@@
    In English legal reasoning tasks, all evaluated LLMs generally exhibit stronger performance compared to their Chinese task. Notably, DeepSeek-R1 and OpenAI o1-preview perform comparably, with each achieving the highest score in three tasks. However, even these advanced LLMs struggle with tasks related to judicial ethics and complex tax calculations.
@@
    Across all legal tasks, the performance of different TTS LLMs models varies significantly. Notably, o1-preview and DeepSeek-R1 outperform other models such as QWQ-32B-Preview and GLM-zero-preview. In the future, further refinement and optimization of these models may be necessary.
  4. 4.
    A detailed error analysis of DeepSeek-R1’s performance on several representative Chinese and English legal tasks, reveals that the model has shortcomings in aspects such as lacking sufficient and updated legal knowledge, struggling to understand legal issues, and factual hallucinations in its thinking process, which offers new insights and directions for enhancing LLMs’ legal reasoning skills.
@@ 2 Related Work
2.1 Test-Time Scaling LLMs
Recently, Test-Time Scaling(TTS) has been an effective way to improve the performance of LLMs. Unlike traditional optimization methods, TTS focuses on optimizing the model’s output during the inference phase through additional computational resources, without requiring modifications to the model’s parameters or structure. This strategy aims to enhance the accuracy and efficiency of the model by dynamically adjusting the reasoning process, particularly when dealing with complex tasks, improving the model’s performance.
A number of works have proposed different methods for enabling LLMs to use test-time scaling for reasoning  Snell et al. (2024); Liu et al. (2025). Existing related works mainly focus on areas such as verifier optimization, proposal distribution refinement, and search algorithms. The verifier optimization involves using process-based verifiers, such as process reward models (PRM)  Lightman et al. (2023); Uesato et al. (2022); Wang et al. (2024) to incrementally evaluate the reasoning steps, which effectively improves the model’s performance in complex reasoning tasks. Additionally, proposal distribution refinement helps LLMs generate more accurate and refined candidate answers during the test phase. This strategy of adjusting the proposal distribution further enhances the model’s reasoning ability, allowing it to iteratively refine its initial outputs and gradually approach the optimal solution. Examples of such methods include STaR and ReSTEM  Singh et al. (2024); Zelikman et al. (2022). These techniques do not require any additional input tokens but instead fine-tune the model to induce an improved proposal distribution. Alternatively, techniques like self-critique  Bai et al. (2022); Du et al. (2023); Madaan et al. (2023); Saunders et al. (2022) enable the model to improve its proposal distribution during test time by instructing it to critique and revise its outputs iteratively. In terms of search algorithms, integrating techniques like Beam Search and Monte Carlo Tree Search (MCTS)  Feng et al. (2024); Trinh et al. (2024) allows for more refined exploration of solution paths, which further enhances the accuracy of reasoning results. A prime example of test-time scaling LLMs is OpenAI’s o1 series models  Openai (2024b), which enhance reasoning capabilities by extending the Chain-of-Thought (CoT) process during the test-time. Besides, Alibaba’s QwQ-32B-Preview  Qwen Team (2024) model showcases the application of test-time computation in mathematical problem-solving. Additionally, the GLM-zero-preview  [^2] https://bigmodel.cn/dev/api/normal-model/glm-zero-preview released by Zhipu AI is another notable example. Recently, by exploring how LLMs can develop reasoning capabilities without supervised data, DeepSeek focuses on the potential for self-evolution through pure RL processes and successfully develops the DeepSeek-R1 model  DeepSeek-AI et al. (2025). Despite having far fewer parameters than o1, DeepSeek-R1 performs on par with OpenAI’s o1-1217 model. Furthermore, DeepSeek-R1 has demonstrated exceptional performance in coding tasks, achieving a 2029 Elo rating on Codeforces, outclassing 96.3% of participants.
However, amid the excitement, there are also concerns, as the TTS LLMs may generate harmful content. If errors occur in legal applications, the consequences can be severe  Weiser (2023). Therefore, it is urgent to test the performance of LLMs in legal contexts. Therefore, this paper aims to explore whether the enhanced abilities of the TTS LLMs effectively transfer to the legal domain and improve their performance on legal reasoning tasks.
@@ 2.2 Legal Reasoning Benchmarks
Understanding the capabilities of LLMs in legal tasks, particularly legal reasoning, is a key focus of research  Blair-Stanek et al. (2023); Trozze et al. (2024). Legal reasoning ability refers to whether LLMs can effectively leverage their legal knowledge and perform the necessary reasoning steps to solve real-world legal tasks  Li et al. (2024a). This requires not only relevant legal knowledge but also the ability to apply that knowledge through reasoning. For instance, in Chinese judicial applications, tasks like legal document generation  Iu and Wong (2023), legal case retrieval  Ma et al. (2021), legal question answering  Hu et al. (2025), and judgment prediction  Gan et al. (2021, 2022); Jiang and Yang (2023); Wei et al. (2025) require that LLMs process precise legal knowledge and complex legal contexts. These tasks often involve specialized components such as judicial interpretation and reasoning.
Many researchers propose various legal reasoning tasks and datasets to assess LLMs’ performance in legal reasoning, covering different legal systems like the civil law system and the common law system  Magesh et al. (2024).  Chlapanis et al. (2024) introduce Legal Argument Reasoning (LAR), a novel task designed to evaluate the legal reasoning capabilities of LLMs.  Deng et al. (2023) develop a syllogistic reasoning dataset for legal judgment analysis, manually curating 11,239 criminal cases that cover 4 criminal elements, 80 charges, and 124 articles.  Joshi et al. (2024) propose the Indian Legal Text Understanding and Reasoning Benchmark (IL-TUR), which includes monolingual (English, Hindi) and multilingual (9 Indian languages) tasks to assess the understanding and reasoning capabilities of Indian legal documents. In addition,  Kang et al. (2023) create a novel corpus focused on Contract Acts in Malaysia and the Australian Social Act for Dependent Children, where ChatGPT is applied using the IRAC method to perform legal analysis.
Several legal task benchmarks are proposed to evaluate whether LLMs meet the high standards required for legal practice, such as Laiw  Dai et al. (2025). In the common law system, LegalBench  Guha et al. (2023) is a widely used legal reasoning benchmark, collaboratively built to consist of 162 tasks across six types of legal reasoning. In the civil law system,  Fei et al. (2024) introduce LawBench, a comprehensive evaluation benchmark that precisely assesses LLMs’ legal capabilities at three cognitive levels: (1) Legal knowledge memorization, (2) Legal knowledge understanding, and (3) Legal knowledge application. Moreover,  Li et al. (2024a) introduce LexEval, a standardized Chinese legal benchmark that focuses not only on LLMs’ ability to apply fundamental legal knowledge but also on examining ethical issues in their application.
However, legal systems vary greatly across countries. The Chinese legal system, rooted in the civil law tradition, differs significantly from the common law systems widely used in the U.S. and the U.K. Therefore, this paper selects a set of existing Chinese and English legal task datasets, along with newly proposed datasets, to comprehensively evaluate the diverse capabilities of LLMs.

### v2

@@ Abstract
Recent advances in test-time scaling of large language models (LLMs), exemplified by DeepSeek-R1 and OpenAI’s o1, show that extending the chain of thought during inference can significantly improve general reasoning performance. However, the impact of this paradigm on legal reasoning remains insufficiently explored. To address this gap, we present the first systematic evaluation of 12 LLMs, including both reasoning-focused and general-purpose models, across 17 Chinese and English legal tasks spanning statutory and case-law traditions. In addition, we curate a bilingual chain-of-thought dataset for legal reasoning through distillation from DeepSeek-R1 and develop Legal-R1, an open-source model specialized for the legal domain. Experimental results show that Legal-R1 delivers competitive performance across diverse tasks. DeepSeek-R1 exhibits clear advantages in Chinese legal reasoning, while OpenAI’s o1 achieves comparable results on English tasks. We further conduct a detailed error analysis, which reveals recurring issues such as outdated legal knowledge, limited capacity for legal interpretation, and susceptibility to factual hallucinations. These findings delineate the main obstacles confronting legal-domain LLMs and suggest promising directions for future research. We release the dataset and model at https://github.com/YinghaoHu/Legal-R1-14B.
Evaluating Test-Time Scaling LLMs for Legal Reasoning: OpenAI o1, DeepSeek-R1, and Beyond
Yinghao Hu 1111Equal contribution., Yaoyao Yu 2,4111Equal contribution., Leilei Gan 3222Corresponding authors., Bin Wei 2,4222Corresponding authors., Kun Kuang 1, Fei Wu 1,5 1College of Computer Science and Technology, Zhejiang University 2Guanghua Law School, Zhejiang University 3School of Software Technology, Zhejiang University 4Law & AI Lab, Zhejiang University 5Shanghai AI Laboratory {huyinghao, yaoyaoyu, leileigan, binwei, kunkuang, wufei}@zju.edu.cn
@@ 1 Introduction
Large language models (LLMs) have recently achieved near-human performance on an increasingly diverse set of benchmarks and application domains Llama3.1; Qwen2.5; openai4o; 2025gemini; anthropic2025claude.
Across several flagship LLM model families, dedicated reasoning variants, such as OpenAI’s o1 openaiLLMs and DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability incorporate an explicit internal deliberation phase before producing a final answer. Fundamentally, these models extend the chain-of-thought (CoT) generated at inference time, thereby allocating increased computational resources per query.
The recent open-sourcing of DeepSeek-R1 further establishes an end-to-end paradigm for training reasoning-centric LLMs. Specifically, deepseekai2025deepseekr1incentivizingreasoningcapability proposes a four-stage pipeline: (i) cold-start pretraining, (ii) reasoning-oriented reinforcement learning (RL), (iii) rejection sampling-based supervised fine-tuning, and (iv) scenario-wide RL. This blueprint has inspired a new wave of test-time computation-intensive models, including QWQ-32B-Preview qwen2024qwq and GLM-zero-preview  [^1] https://bigmodel.cn/dev/api/normal-model/glm-zero-preview, which similarly extend reasoning traces, trading off computational cost for improved inference accuracy.
Figure 1: Overview of Work. The figure presents the 12 evaluated models together with representative task types and their data sources.
Contemporaneous work explores inference-time search strategies and training signals, such as Process Reward Models lightman2023letsverifystepstep; uesato2022solvingmathwordproblems; wang-etal-2024-math, self-corrective RL schemes kumar2024traininglanguagemodelsselfcorrect, and Monte Carlo Tree Search (MCTS) and beam search variants feng2024alphazeroliketreesearchguidelarge; Trinh53097. While these approaches have not yet matched the reported performance of o1 openaiLLMs and DeepSeek-R1, they nonetheless offer valuable insights for advancing the capabilities of reasoning-focused LLMs.
Figure 2: Overall Performance of LLMs on Chinese and English Legal Tasks. The figure shows the performance of representative LLMs on Chinese and English legal tasks. Inference models such as DeepSeek-R1 and o1-preview outperform traditional LLMs, while our model Legal-R1 achieves competitive performance.
While the reasoning capabilities of LLMs have improved substantially in recent years, it would be premature to assume that such progress necessarily translates into strong performance on legal tasks. Legal reasoning imposes two simultaneous and demanding requirements: (i) the accurate synthesis of relevant statutes and case knowledge, and (ii) the rigorous application of this knowledge to novel and often complex fact patterns. Consequently, it remains uncertain whether models that perform well on general-purpose reasoning benchmarks can satisfy the domain-specific demands of legal reasoning. Although prior work has examined GPT-4 on specific legal tasks—such as legal text annotation Savelka2023, explaining legislative terminology savelka2023explaininglegalconceptsaugmented, and thematic analysis in empirical legal studies drápal2023usinglargelanguagemodels—these studies target narrow applications and assess models not purpose-built for reasoning. Consequently, a systematic evaluation of LLMs’ legal reasoning across tasks spanning both statutory and case-law systems is still lacking.
To address this gap, we (i) present the first systematic evaluation of 17 legal reasoning tasks — seven in English and ten in Chinese — covering both test-time scaled and general-purpose LLMs; and (ii) construct a bilingual legal reasoning dataset using rejection sampling. Using this dataset, we progressively fine-tune DeepSeek-R1-Distill-Qwen-14B via supervised learning, resulting in Legal-R1, a domain-specific model with enhanced performance on legal tasks. Finally, we analyze errors across representative Chinese- and English-language tasks, identifying key challenges and future directions for improving legal reasoning in LLMs.
Our contributions can be summarized as follows:
@@
    Among the evaluated models, DeepSeek-R1 demonstrates superior performance in Chinese legal reasoning tasks. In English settings, both models perform similarly, achieving top results across several tasks. Nevertheless, even the strongest models continue to struggle with advanced reasoning tasks, such as those involving judicial ethics and complex tax calculations.
@@
    We introduce Legal-R1, developed through a progressive supervised fine-tuning strategy. It outperforms baseline models on the majority of Chinese and English legal tasks and exceeds DeepSeek-R1 on key tasks such as LC and IAPE, establishing a new standard for legal reasoning.
@@
    Our error analysis on representative Chinese and English legal tasks reveals key weaknesses, including outdated knowledge, limited legal understanding, and factual hallucinations. These results point to important directions for enhancing legal reasoning in LLMs.
@@ 2.1 Legal Reasoning Benchmarks
Understanding the capabilities of LLMs in legal tasks, particularly legal reasoning, is a key focus of research  blairstanek2023gpt3performstatutoryreasoning; trozze2024large, especially in tasks such as legal document generation  Iu2023ChatGPTBO, question answering  hu2025fine, and judgment prediction  gan2021judgment; gan2022exploiting; JiangLegal2023; wei2025llms; yuan2024can; yuan2026multi. To facilitate legal reasoning evaluation, researchers have developed a diverse range of legal benchmarks, including LAR-ECHR chlapanis-etal-2024-lar and IL-TUR joshi-etal-2024-il. In addition, comprehensive benchmark suites such as LegalBench LEGALBENCH for common-law tasks, LawBench fei-etal-2024-lawbench for civil-law evaluation, LexEval li2024lexevalcomprehensivechineselegal for Chinese legal texts with ethical considerations, and Laiw dai-etal-2025-laiw, which emphasizes practice-oriented criteria, have been introduced.
However, legal systems differ across jurisdictions. Therefore, we construct a set of legal reasoning datasets covering both Chinese and U.S. legal systems to comprehensively evaluate the legal reasoning capabilities of current LLMs.
2.2 Test-Time Scaling
TTS has emerged as a powerful technique to boost the reasoning capabilities of LLMs during inference, without altering their underlying parameters or architecture. This paradigm has been adopted by several prominent models, including OpenAI’s o1 series openaiLLMs, Alibaba’s QwQ-32B-Preview qwen2024qwq, Zhipu AI’s GLM-zero-preview, and DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability. Several methods have been proposed to enable LLMs to leverage test-time scaling for enhanced reasoning. Verifier optimization, for instance, through process reward models lightman2023letsverifystepstep; uesato2022solvingmathwordproblems; wang-etal-2024-math, facilitates the incremental evaluation of reasoning steps, thereby boosting performance on complex tasks. Methods like STaRzelikman2022starbootstrappingreasoningreasoning and ReST singh2024humandatascalingselftraining refine proposal distributions by fine-tuning models to generate more accurate answers without adding extra tokens. Self-critique techniques  bai2022constitutionalaiharmlessnessai; du2023improvingfactualityreasoninglanguage; madaan2023selfrefineiterativerefinementselffeedback; saunders2022selfcritiquingmodelsassistinghuman allow the model to iteratively refine its outputs. Search algorithms like Beam Search and Monte Carlo Tree Searchfeng2024alphazeroliketreesearchguidelarge; Trinh53097 further enhance exploration and solution accuracy. Despite their promise, the effectiveness of TTS-enhanced LLMs in legal tasks remains underexplored. This paper investigates whether their improved reasoning capabilities can transfer to the legal domain.